Solve by the elimination method. x - y = -13 and 6x - 8y = -86
I need this today please. Thank you.

Answer by macston(5194)

6x - 8y = -86
x - y = -13 Multiply this equation by -6
6x - 8y = -86
-6x + 6y = 78 and add the two equations
__________
-2y=-8
y=4 ANSWER 1 y=4 solve either equation for x
x-y=-13
x-4=-13
x=-9 ANSWER 2 x=-9
CHECK (use the equation you did not use to solve for x
6x-8y=-86
6(-9)-8(4)=-86
-54-32=-86
-86=-86